Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies

by leumas » Sat Sep 10, 2011 2:25 am
prateek_guy2004 wrote:
[email protected] wrote:The probability of having a girl in all 3 cases is :
=> 1/2*1/2*1/2= 1/8

The probability of NOT having a boy in all 3 cases is :
=> 1/2*1/2*1/2= 1/8

Therefore the probability of having a girl and not a boy will be;

=> 1/8 +1/8
=> 2/8
=> 1/4

Ans : 1/4
This is a easy approach.....
I feel this is not answering the question correctly. Either all boy or all girl. It answers all girls and not a boy.

Somebody clarify.
Join the discussion

by JS_2 » Mon Sep 12, 2011 10:03 pm
1/4
2 favourable outcomes: either all boys or all girls
All possible outcomes: (either boy or girl=2)*(either boy or girl=2)*(either boy or girl=2)

i.e, 2/2*2*2= 2/8=1/4.
Join the discussion

by prashant misra » Thu Sep 15, 2011 7:00 am
thank you stuart sir for explaining different approach towards the question.The answer is D
Join the discussion

by vini1612 » Tue Sep 20, 2011 8:32 am
niazsna786 wrote:The question says, the probability of all the 3 children be same gender

The children can be all 3 girls or all 3 boys.

The probability of getting a boy = 1/2
The probability of getting a girls = 1/2


Probability of getting 3 girls = 1/2 * 1/2 * 1/2 = 1/8
Probability of getting 3 boys= 1/2 * 1/2 * 1/2 = 1/8

What is asked is probability of getting of same gender i.e. either all 3 girls or all 3 boys i.e. 1/8 + 1/8 = 1/4
E is the answer
Thanks! this made me understand the solution easily!
Join the discussion

by parul9 » Fri Oct 07, 2011 11:56 pm
E - 1/4!
Join the discussion

by nsamkari » Sat Oct 22, 2011 11:35 pm
The answer is 1/4
Join the discussion

by Deependra1 » Wed Oct 26, 2011 4:40 am
Answer: A
Join the discussion

by Sharma_Gaurav » Mon Jan 09, 2012 3:38 pm
straight E is the answer
Join the discussion

by mparakala » Wed Nov 07, 2012 8:01 am
E

NS: 2/8 = 1/4
Join the discussion

by rajeshsinghgmat » Mon Jan 28, 2013 12:07 am
sEE the ansEr.

P(E)= (1/2)*(1/2)*(1/2) + (1/2)*(1/2)*(1/2)
= (1/2)^3 + (1/2)^3
= 1/8 +1/8
= 2/8
= 1/4
Join the discussion

by Mathsbuddy » Thu Nov 28, 2013 12:35 am
P(1st child male or female) = 1
P(2nd child matching previous) = 1/2
P(3rd child matching previous) = 1/2

Total probability = 1 * 1/2 * 1/2 = 1/4
Join the discussion

by lulufrenchie » Wed Apr 09, 2014 11:21 pm
Hi all,

I draw a probability tree to answer this question and found the right answer.

But is it a waist of time ?
Join the discussion

by Brent@GMATPrepNow » Thu Apr 10, 2014 8:08 am
lulufrenchie wrote:Hi all,

I draw a probability tree to answer this question and found the right answer.

But is it a waist of time ?
A probability tree doesn't take long AND it provides an easy way to SEE all of the possible outcomes:
Image

Of the 8 possible outcomes, let's highlight those where all 3 children are the same gender.
Image

There are 2 such outcomes, so the probability = 2/8 = [spoiler]1/4 = E[/spoiler]

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by bml1105 » Tue May 20, 2014 6:46 pm
I know this is an old thread, but what's the difference between the original question and this question:

"A couple decides to have 4 children. If they succeed in having 4 children and each child is equally likely to be a boy or a girl, what is the probability that they will have exactly 2 girls and 2 boys?"

(A) 3/8
(B) 1/4
(C) 3/16
(D) 1/8
(E) 1/16


In the problem I posted, the solution is to use permutation/combination instead of multiplying the fractions
Join the discussion

by [email protected] » Tue May 20, 2014 11:12 pm
Hi bml1105,

Your question can also be solved in a variety of ways. Since there are 4 children and each is either a boy or a girl, then you could draw a probability tree (there are 2^4 = 16 possibilities) or you could use permutations/combinations or you could use probability math. Most questions of this type can be solved in more than 1 way, so the measure of your approach is if it gets you the correct answer AND is relatively fast to use.

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion